# Grounding Long-Context Reasoning with Contextual Normalization for Retrieval-Augmented Generation

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) has become an essential approach for extending the reasoning and knowledge capacity of large language models (LLMs). While prior research has primarily focused on retrieval quality and prompting strategies, the influence of how the retrieved documents are framed, i.e., context format, remains underexplored. We show that seemingly superficial choices, such as delimiters or structural markers in key–value extraction, can induce substantial shifts in accuracy and stability, even when semantic content is identical. To systematically investigate this effect, we design controlled experiments that vary context density, delimiter styles, and positional placement, revealing the underlying factors that govern performance differences. Building on these insights, we introduce Contextual Normalization, a lightweight strategy that adaptively standardizes context representations before generation. Extensive experiments on both controlled and real-world RAG benchmarks across diverse settings demonstrate that the proposed strategy consistently improves robustness to order variation and strengthens long-context utilization. These findings underscore that reliable RAG depends not only on retrieving the right content, but also on how that content is presented, offering both new empirical evidence and a practical technique for better long-context reasoning.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has emerged as a foundational paradigm for enabling large language models (LLMs) to scale to knowledge-intensive tasks by conditioning generation on external documents retrieved from large corpora (Lewis et al., 2020b; Borgeaud et al., 2022). In a standard pipeline, a retriever first identifies potentially relevant texts for a given query, and these documents are then concatenated into a prompt for the LLM. With the advent of long-context LLMs that can process tens of thousands of tokens (Xiao et al., 2024; Xu et al., 2024), the opportunities for complex reasoning over vast information spaces have been unlocked, making RAG increasingly central to real-world applications such as open-domain QA and scientific literature analysis.
@@
While long-context extensions enable RAG systems to scale to much larger evidence pools, they also introduce new challenges. Recent study (Leng et al., 2024) highlights these limitations by systematically varying context length, from 2K up to 128K tokens across dozens of models, and documenting consistent failure modes when contexts become too long or unwieldy. With the number of retrieved chunks increasing, LLMs face amplified retrieval noise, redundancy across overlapping documents, and dilution of truly relevant evidence. These issues often make it harder for LLMs to distinguish signal from distraction, leading to unstable reasoning and degraded accuracy. Moreover, positional biases (Liu et al., 2024a; Zhang et al., 2024) further interact with these challenges: LLMs tend to over-attend to the beginning or end of a prompt, leaving evidence buried in the middle underutilized. Together, these factors expose a fundamental brittleness in long-context RAG that limits its reliability in real-world deployments.
@@
To mitigate these limitations, a growing line of research has explored strategies to improve long-context RAG performance. One representative approach is the prompt optimization (Liu et al., 2024b), where multiple permutations of retrieved chunks are scored, and the prompt yielding the highest likelihood is selected for answering. Another direction relies on synthetic supervision: An et al. (2024) proposes constructing curated datasets where answers depend on specific chunks within extended inputs, encouraging LLMs to develop position-invariant reasoning strategies. While effective in controlled settings, such methods face scalability issues, as generating chunk-level annotations is costly and risks positional overfitting. Architectural modifications, such as redesigned positional encodings (Zhang et al., 2024), offer more fundamental solutions but require non-trivial changes to model internals. Complementary work (Vladika & Matthes, 2025) provides further evidence that context size, snippet count, and model architecture interact in subtle ways, jointly shaping robustness and accuracy.
@@
To benchmark long-context reasoning in a way that isolates potential factors from prior knowledge, Liu et al. (2024a) propose the key–value extraction task, where LLMs must retrieve the correct value for a given key from a synthetic context. Inspired by this controlled setup, we extend the analysis and uncover a striking finding. As illustrated in Figure 1, even when semantics and input length are held constant, altering the surface format of key–value pairs, for instance, representing them as UUIDs, plain texts, or switching delimiters such as “-” versus “&”, leads to substantial performance differences. This finding highlights that the presentation format of context, beyond its
@@ 2 How Context Format Grounds
The performance of RAG systems in long-context scenarios is profoundly influenced by the effective integration of retrieved information (Borgeaud et al., 2022; Karpukhin et al., 2020). While previous work (Asai et al., 2023) has primarily focused on the quantity and relevance of retrieved documents, we posit that the internal format of this information, more specifically, how context content in each chunk is structured, plays a critical role in grounding the model’s generation. To investigate this hypothesis, we design a set of experiments centered on the key-value extraction task (Liu et al., 2024a), a canonical challenge for LLMs requiring precise information retrieval from a given context. The goal of the task is to retrieve the value of a specific key from a long JSON object. More details are provided in Appendix A.
Context Formats. To systematically investigate how context format grounds LLM’s generation, we propose the following context formats in the experiments: UUID, Plain Text, and Modified UUID. Each format varies only in its use of structured identifiers, allowing us to analyze the model performance to metadata and special characters. The Universally Unique Identifiers (UUIDs) is a 128-bit number used to uniquely identify information in computer systems, which is utilized in the standard key-value retrieval task (Liu et al., 2024a). They are typically represented as a 32-digit hexadecimal number, displayed in five groups separated by hyphens. In Plain Text, all structured identifiers are removed. Both key and value are flattened into a continuous text string. Modified UUID introduces a subtle change to the original UUID by replacing the hyphen (-) with a different delimiter (&). This simple substitution allows us to probe the sensitivity of LLMs to variations in structured data, revealing how the model processes the context.
Settings. We design a controlled experiment with 500 samples to evaluate the impact of context format on long-context RAG. We permute the position of a “gold” key within a long context, then measure the performance of three LLMs: LLaMA-2-7B (Touvron et al., 2023), LLaMA-2-7B-Chat (Touvron et al., 2023), and Qwen2.5-1.5B (Yang et al., 2024). The experiments are conducted with two context configurations to test different densities: low-density (40 contexts with 32 characters each) and high-density (10 contexts with 128 characters each). For each setting, we record two key metrics: the Overall Averaged Accuracy (OAA) across all positions, which measures robustness to all gold key positions, and the Optimal Positioned Accuracy (OPA), which captures the model’s best-case performance under ideal position of gold key.
@@ 3.1 Tokenization
We first look into the impact of tokenization on this key-value extraction task, focusing specifically on how delimiter choices interact with the tokenizer internals. For LLMs such as Qwen2.5, which use a SentencePiece-based tokenizer (Kudo & Richardson, 2018), delimiter characters such as ‘-’, ‘:’, ‘&’, ‘_’, and ‘+’ affect the token count of the input string significantly. To this end, we use 200
@@
On the Role of Training Data. To further probe why different context formats lead to distinct attention allocation patterns, we attempt to trace the effect back to the training data. With Stanford-Alpaca-7B (Taori et al., 2023), we sort tokens in its fine-tuning corpus by frequency of occurrence, and then reconstruct QA contexts where original tokens are replaced with either the most frequent or least frequent tokens. This design tests whether exposure frequency in fine-tuning data influences how attention is distributed across contexts. However, the results do not show a clear relationship between token frequency and LLM performance or attention allocation. This indicates that the grounding mechanism behind context-format sensitivity is more complex than simple token frequency statistics, likely shaped by deeper patterns acquired during both pretraining and fine-tuning. We provide the details of the experiment in Appendix B.
@@ 5.1 Controlled Long-Context RAG Settings
In this case, we propose a controlled test using a permuted version of NQ-Open to evaluate both the robustness to order variation and long-context reasoning capacity of LLMs. First, we randomly sample 500 questions from NQ-Open (Liu et al., 2024a). For each question, one gold (relevant) document is identified and mixed with 9 distractors, each containing about 100–300 tokens. We then construct 10 input permutations by placing the gold document at each possible position while shuffling the remaining distractors. The ratio $p$ in C-Normis fixed at $p=0.5$ with 8 samples used for selecting the best delimiters.
@@
Models. We adopt several LLMs for evaluation, including LLaMA-2-7B (pretrained context length 4K) (Touvron et al., 2023), LLaMA-2-7B-Chat (4K), Qwen2.5-1.5B (128K) (Yang et al., 2024), and Qwen2.5-1.5B-Instruct (128K). For base models (e.g., LLaMA-2-7B and Qwen2.5-1.5B), we use unaligned prompts directly following the QA format. For instruction-tuned models (e.g., LLaMA-2-7B-Chat and Qwen2.5-1.5B-Instruct), we adopt aligned prompts that match their chat/instruction interfaces. All generations are performed with temperature fixed at 0 to ensure deterministic outputs and eliminate randomness from sampling.
@@ 5.2 Real-World RAG Settings
To evaluate the real-world utility of C-Norm, we adopt LongBench-v2 (Bai et al., 2024), a benchmark targeting long-context reasoning across diverse tasks. It contains 503 multiple-choice questions drawn from six categories, including single-document QA, multi-document QA, long in-context learning, dialogue history understanding, codebase comprehension, and structured data understanding. It covers both textual and semi-structured formats. Each question is paired with a long context ranging from 8K to over 2M words, with most falling under 128K, making it ideal for testing long-context generalization. We evaluate under two settings:
@@ 6 Related Work
Retrieval Augmented Generation (RAG) has been widely adopted to improve language models’ performance on knowledge-intensive tasks (Borgeaud et al., 2022; Lewis et al., 2020b; Karpukhin et al., 2020). Traditional RAG pipelines usually manage short context windows, typically involving tasks with concise and immediately relevant contexts (Lewis et al., 2020a). While effective for short and well-contained queries, the systems face substantial limitations when scaling to more complex or open-ended tasks (Jeong et al., 2024). Many real-world questions require integrating dispersed evidence from multiple documents or reasoning over lengthy documents such as academic articles, legal cases, or multi-turn dialogues. Standard pipelines that retrieve and concatenate only a few short passages (typically 100–300 tokens each) often suffer from information fragmentation or omission of critical context (Li et al., 2024b; Hsieh et al., 2024). Furthermore, fixed-length context windows in most pretrained LLMs (e.g., 2K–4K tokens) severely limit the amount of retrievable evidence considered simultaneously. These bottlenecks have prompted shifts toward long-context RAG setups, aiming to leverage larger contexts and improved retrieval for open-domain QA (Asai et al., 2023; Lee et al., 2019; Nakano et al., 2021; Li et al., 2025), multi-hop reasoning (Zhong et al., 2023; Ho et al., 2020), and complex document understanding (Dua et al., 2019; Li et al., 2024a).
Long-Context RAG. Recent studies (Liu et al., 2024a; Zhang et al., 2024; An et al., 2024; Liu et al., 2024b) have revealed critical limitations in how large language models utilize long-context inputs in RAG. Simply appending more retrieved text does not guarantee improved performance, potentially causing degradation due to positional biases, information dilution, and the “lost-in-the-middle” phenomenon (Liu et al., 2024a). Models often favor content at the beginning or end of the prompt, neglecting relevant information buried in the middle. This results in significant performance variance depending on the order of retrieved documents, even if the overall content remains unchanged (Liu et al., 2024b; Zhang et al., 2024; An et al., 2024). Thus, the effectiveness of long-context RAG is influenced not only by the amount of available information but also by how it is ordered and integrated, motivating a deeper empirical analysis of context-order effects on LLM performance.
@@ Appendix B Frequency-Controlled Token Replacement
Settings. We evaluate the robustness and capacity of LLM on 100 samples from the NQ-Open dataset (Liu et al., 2024a). Each sample is paired with 6 retrieved documents, each containing approximately 100–300 tokens. To simulate long-context reasoning, we permute the position of the gold document across all possible positions. For token replacement, we sort tokens in the Alpaca fine-tuning data by frequency of occurrence and define replacement groups corresponding to the top-$k\%$ most frequent tokens and bottom-$k\%$ least frequent tokens ($k=1,5,10$). Replacement is enforced by prompting the model to rewrite retrieved passages using only tokens from the allowed set, according to the following instruction:

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) has become an essential approach for extending the reasoning and knowledge capacity of large language models (LLMs). While prior research has primarily focused on retrieval quality and prompting strategies, the influence of how the retrieved documents are framed, i.e., context format, remains underexplored. We show that seemingly superficial choices, such as delimiters or structural markers in key-value extraction, can induce substantial shifts in accuracy and stability, even when semantic content is identical. To systematically investigate this effect, we design controlled experiments that vary context density, delimiter styles, and positional placement, revealing the underlying factors that govern performance differences. Building on these insights, we introduce Contextual Normalization, a lightweight strategy that adaptively standardizes context representations before generation. Extensive experiments on both controlled and real-world RAG benchmarks across diverse settings demonstrate that the proposed strategy consistently improves robustness to order variation and strengthens long-context utilization. These findings underscore that reliable RAG depends not only on retrieving the right content, but also on how that content is presented, offering both new empirical evidence and a practical technique for better long-context reasoning.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has emerged as a foundational paradigm for enabling large language models (LLMs) to scale to knowledge-intensive tasks by conditioning generation on external documents retrieved from large corpora (lewis2020retrieval; borgeaud2022improving). In a standard pipeline, a retriever first identifies potentially relevant texts for a given query, and these documents are then concatenated into a prompt for the LLM. With the advent of long-context LLMs that can process tens of thousands of tokens (DBLP:conf/iclr/XiaoTCHL24; DBLP:conf/iclr/0008PWM0LSBSC24), the opportunities for complex reasoning over vast information spaces have been unlocked, making RAG increasingly central to real-world applications such as open-domain QA and scientific literature analysis.
@@
While long-context extensions enable RAG systems to scale to much larger evidence pools, they also introduce new challenges. Recent study (leng2024long) highlights these limitations by systematically varying context length, from 2K up to 128K tokens across dozens of models, and documenting consistent failure modes when contexts become too long or unwieldy. With the number of retrieved chunks increasing, LLMs face amplified retrieval noise, redundancy across overlapping documents, and dilution of truly relevant evidence. These issues often make it harder for LLMs to distinguish signal from distraction, leading to unstable reasoning and degraded accuracy. Moreover, positional biases (DBLP:journals/tacl/LiuLHPBPL24; zhangfound) further interact with these challenges: LLMs tend to over-attend to the beginning or end of a prompt, leaving evidence buried in the middle underutilized. Together, these factors expose a fundamental brittleness in long-context RAG that limits its reliability in real-world deployments.
@@
To mitigate these limitations, a growing line of research has explored strategies to improve long-context RAG performance. One representative approach is the prompt optimization (liu2024likelihood), where multiple permutations of retrieved chunks are scored, and the prompt yielding the highest likelihood is selected for answering. Another direction relies on synthetic supervision: an2024make proposes constructing curated datasets where answers depend on specific chunks within extended inputs, encouraging LLMs to develop position-invariant reasoning strategies. While effective in controlled settings, such methods face scalability issues, as generating chunk-level annotations is costly and risks positional overfitting. Architectural modifications, such as redesigned positional encodings (zhangfound), offer more fundamental solutions but require non-trivial changes to model internals. Complementary work (vladika2025influence) provides further evidence that context size, snippet count, and model architecture interact in subtle ways, jointly shaping robustness and accuracy.
@@
To benchmark long-context reasoning in a way that isolates potential factors from prior knowledge, DBLP:journals/tacl/LiuLHPBPL24 propose the key–value extraction task, where LLMs must retrieve the correct value for a given key from a synthetic context. Inspired by this controlled setup, we extend the analysis and uncover a striking finding. As illustrated in Figure 1, even when semantics and input length are held constant, altering the surface format of key–value pairs, for instance, representing them as UUIDs, plain texts, or switching delimiters such as “-” versus “&”, leads to substantial performance differences. This finding highlights that the presentation format of context, beyond its
@@ 2 How Context Format Grounds
The performance of RAG systems in long-context scenarios is profoundly influenced by the effective integration of retrieved information (borgeaud2022improving; karpukhin2020dense). While previous work (DBLP:conf/acl/AsaiSL0I0HY23) has primarily focused on the quantity and relevance of retrieved documents, we posit that the internal format of this information, more specifically, how context content in each chunk is structured, plays a critical role in grounding the model’s generation. To investigate this hypothesis, we design a set of experiments centered on the key-value extraction task (DBLP:journals/tacl/LiuLHPBPL24), a canonical challenge for LLMs requiring precise information retrieval from a given context. The goal of the task is to retrieve the value of a specific key from a long JSON object. More details are provided in Appendix A.
Context Formats. To systematically investigate how context format grounds LLM’s generation, we propose the following context formats in the experiments: UUID, Plain Text, and Modified UUID. Each format varies only in its use of structured identifiers, allowing us to analyze the model performance to metadata and special characters. The Universally Unique Identifiers (UUIDs) is a 128-bit number used to uniquely identify information in computer systems, which is utilized in the standard key-value retrieval task (DBLP:journals/tacl/LiuLHPBPL24). They are typically represented as a 32-digit hexadecimal number, displayed in five groups separated by hyphens. In Plain Text, all structured identifiers are removed. Both key and value are flattened into a continuous text string. Modified UUID introduces a subtle change to the original UUID by replacing the hyphen (-) with a different delimiter (&). This simple substitution allows us to probe the sensitivity of LLMs to variations in structured data, revealing how the model processes the context.
Settings. We design a controlled experiment with 500 samples to evaluate the impact of context format on long-context RAG. We permute the position of a “gold” key within a long context, then measure the performance of three LLMs: LLaMA-2-7B (llama), LLaMA-2-7B-Chat (llama), and Qwen2.5-1.5B (qwen2.5). The experiments are conducted with two context configurations to test different densities: low-density (40 contexts with 32 characters each) and high-density (10 contexts with 128 characters each). For each setting, we record two key metrics: the Overall Averaged Accuracy (OAA) across all positions, which measures robustness to all gold key positions, and the Optimal Positioned Accuracy (OPA), which captures the model’s best-case performance under ideal position of gold key.
@@ 3.1 Tokenization
We first look into the impact of tokenization on this key-value extraction task, focusing specifically on how delimiter choices interact with the tokenizer internals. For LLMs such as Qwen2.5, which use a SentencePiece-based tokenizer (DBLP:conf/emnlp/KudoR18), delimiter characters such as ‘-’, ‘:’, ‘&’, ‘_’, and ‘+’ affect the token count of the input string significantly. To this end, we use 200
@@
On the Role of Training Data. To further probe why different context formats lead to distinct attention allocation patterns, we attempt to trace the effect back to the training data. With Stanford-Alpaca-7B (alpaca), we sort tokens in its fine-tuning corpus by frequency of occurrence, and then reconstruct QA contexts where original tokens are replaced with either the most frequent or least frequent tokens. This design tests whether exposure frequency in fine-tuning data influences how attention is distributed across contexts. However, the results do not show a clear relationship between token frequency and LLM performance or attention allocation. This indicates that the grounding mechanism behind context-format sensitivity is more complex than simple token frequency statistics, likely shaped by deeper patterns acquired during both pretraining and fine-tuning. We provide the details of the experiment in Appendix B.
@@ 5.1 Controlled Long-Context RAG Settings
In this case, we propose a controlled test using a permuted version of NQ-Open to evaluate both the robustness to order variation and long-context reasoning capacity of LLMs. First, we randomly sample 500 questions from NQ-Open (DBLP:journals/tacl/LiuLHPBPL24). For each question, one gold (relevant) document is identified and mixed with 9 distractors, each containing about 100–300 tokens. We then construct 10 input permutations by placing the gold document at each possible position while shuffling the remaining distractors. The ratio $p$ in C-Normis fixed at $p=0.5$ with 8 samples used for selecting the best delimiters.
@@
Models. We adopt several LLMs for evaluation, including LLaMA-2-7B (pretrained context length 4K) (llama), LLaMA-2-7B-Chat (4K), Qwen2.5-1.5B (128K) (qwen2.5), and Qwen2.5-1.5B-Instruct (128K). For base models (e.g., LLaMA-2-7B and Qwen2.5-1.5B), we use unaligned prompts directly following the QA format. For instruction-tuned models (e.g., LLaMA-2-7B-Chat and Qwen2.5-1.5B-Instruct), we adopt aligned prompts that match their chat/instruction interfaces. All generations are performed with temperature fixed at 0 to ensure deterministic outputs and eliminate randomness from sampling.
@@ 5.2 Real-World RAG Settings
To evaluate the real-world utility of C-Norm, we adopt LongBench-v2 (DBLP:journals/corr/abs-2412-15204), a benchmark targeting long-context reasoning across diverse tasks. It contains 503 multiple-choice questions drawn from six categories, including single-document QA, multi-document QA, long in-context learning, dialogue history understanding, codebase comprehension, and structured data understanding. It covers both textual and semi-structured formats. Each question is paired with a long context ranging from 8K to over 2M words, with most falling under 128K, making it ideal for testing long-context generalization. We evaluate under two settings:
@@ 6 Related Work
Retrieval Augmented Generation (RAG) has been widely adopted to improve language models’ performance on knowledge-intensive tasks (borgeaud2022improving; lewis2020retrieval; karpukhin2020dense). Traditional RAG pipelines usually manage short context windows, typically involving tasks with concise and immediately relevant contexts (DBLP:conf/nips/LewisPPPKGKLYR020). While effective for short and well-contained queries, the systems face substantial limitations when scaling to more complex or open-ended tasks (DBLP:conf/naacl/JeongBCHP24). Many real-world questions require integrating dispersed evidence from multiple documents or reasoning over lengthy documents such as academic articles, legal cases, or multi-turn dialogues. Standard pipelines that retrieve and concatenate only a few short passages (typically 100–300 tokens each) often suffer from information fragmentation or omission of critical context (li2024retrieval; DBLP:journals/corr/abs-2404-06654). Furthermore, fixed-length context windows in most pretrained LLMs (e.g., 2K–4K tokens) severely limit the amount of retrievable evidence considered simultaneously. These bottlenecks have prompted shifts toward long-context RAG setups, aiming to leverage larger contexts and improved retrieval for open-domain QA (DBLP:conf/acl/AsaiSL0I0HY23; DBLP:conf/acl/LeeCT19; DBLP:journals/corr/abs-2112-09332; li2025towards), multi-hop reasoning (DBLP:conf/emnlp/ZhongWMPC23; DBLP:conf/coling/HoNSA20), and complex document understanding (DBLP:conf/naacl/DuaWDSS019; DBLP:journals/corr/abs-2410-03227).
Long-Context RAG. Recent studies (DBLP:journals/tacl/LiuLHPBPL24; zhangfound; an2024make; liu2024likelihood) have revealed critical limitations in how large language models utilize long-context inputs in RAG. Simply appending more retrieved text does not guarantee improved performance, potentially causing degradation due to positional biases, information dilution, and the “lost-in-the-middle” phenomenon (DBLP:journals/tacl/LiuLHPBPL24). Models often favor content at the beginning or end of the prompt, neglecting relevant information buried in the middle. This results in significant performance variance depending on the order of retrieved documents, even if the overall content remains unchanged (liu2024likelihood; zhangfound; an2024make). Thus, the effectiveness of long-context RAG is influenced not only by the amount of available information but also by how it is ordered and integrated, motivating a deeper empirical analysis of context-order effects on LLM performance.
@@ Appendix B Frequency-Controlled Token Replacement
Settings. We evaluate the robustness and capacity of LLM on 100 samples from the NQ-Open dataset (DBLP:journals/tacl/LiuLHPBPL24). Each sample is paired with 6 retrieved documents, each containing approximately 100–300 tokens. To simulate long-context reasoning, we permute the position of the gold document across all possible positions. For token replacement, we sort tokens in the Alpaca fine-tuning data by frequency of occurrence and define replacement groups corresponding to the top-$k\%$ most frequent tokens and bottom-$k\%$ least frequent tokens ($k=1,5,10$). Replacement is enforced by prompting the model to rewrite retrieved passages using only tokens from the allowed set, according to the following instruction: